# On dissipation operators of Quantum Optics

### v3

@@ Abstract
We study properties of two basic dissipation operators of Quantum Optics in the framework of the theory of the Hilbert space of the Hilbert-Schmidt operators. We prove that one of the operators is symmetric and nonpositive, while another is neither symmetric nor nonpositive.
@@ 2 Dissipation operators and main results
$\rho=\sum_{n,n^{\prime}}\,\sum_{s,s^{\prime}=s_{\pm}}\rho_{n,n^{\prime}}|n,s\rangle\otimes\langle n^{\prime},s^{\prime}|.$ (2.2)
@@
$\rho_{n,n^{\prime}}=\langle n|\rho|n^{\prime}\rangle,$ (2.3)
@@
$\|\rho\|_{\rm HS}^{2}=\mathop{\rm tr\,}\nolimits[\rho^{2}]=\sum_{n,n^{\prime}=0}^{\infty}\mathop{\rm tr\,}\nolimits\rho^{2}_{n,n^{\prime}}<\infty.$ (2.4)
Definition $2.$
@@
i) The dissipation operator $D$ is defined on the domain ${\cal D}\subset{\rm HS}$ and it is nonpositive and symmetric:
$\langle\rho,D\rho\rangle_{\rm HS}\leq 0,\qquad\langle\rho,D\rho\rangle_{\rm HS}=\langle D\rho,\rho\rangle_{\rm HS},\qquad\rho\in{\cal D}.$ (2.6)
ii) The operator ${\Delta}:{\cal D}\to{\cal D}$ is neither symmetric nor nonpositive.
iii) The operator $D:{\cal D}\to{\cal D}$ is injective.
@@
Let us comment on related previous results. The dissipation operators $\Delta$} and $D$ were introduced analysing quantum spontaneous emission in [1][3] and [4, 5, 10, 12, 13].
In [7] and [8], we have demonstrated the role of classical dissipation in adiabatic single-frequency asymptotics for solutions to the semiclassical Maxwell–Bloch equations.
@@ 3 Nonpositivity and symmetry
Here we prove Theorem 2.3 i) and ii).
i) First, let us prove the nonpositivity of $D$. For $\rho\in{\cal D}$,
$\langle\rho,D\rho\rangle_{\rm HS}$ $=$ $\mathop{\rm tr\,}\nolimits\Big{(}\rho D\rho\Big{)}=\mathop{\rm tr\,}\nolimits\Big{(}\rho\big{[}a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho(t)-\frac{1}{2}\rho a^{\dagger}a+a^{\dagger}\rho a-\frac{1}{2}aa^{\dagger}\rho-\frac{1}{2}\rho aa^{\dagger}\big{]}\Big{)}$ (3.1)
$=$ $\mathop{\rm tr\,}\nolimits\Big{(}\rho a\rho a^{\dagger}-\rho a^{\dagger}a\rho+\rho a^{\dagger}\rho a-\rho aa^{\dagger}\rho\Big{)}=\mathop{\rm tr\,}\nolimits\Big{(}2\rho a\rho a^{\dagger}-a^{\dagger}a\rho^{2}-aa^{\dagger}\rho^{2}\Big{)}.$
@@
Now we use the fact that $\rho$ is a finite rank Hermitian operator (2.2). Then all operators $\rho a\rho a^{\dagger}$, $a^{\dagger}a\rho^{2}$ and $aa^{\dagger}\rho^{2}$ have only finite number of nonzero entries (2.3), so their traces are well defined. Moreover, $\rho$ admits a finite spectral resolution $\rho=(\rho_{i}\delta_{ij})$ in the orthonormal basis of its eigenvectors from $X_{0}$:
$\rho=\sum_{i=1}^{\infty}\rho_{i}\,e_{i}\!\otimes\!e_{i},\qquad e_{i}\in X_{0},$ (3.2)
@@
where only finite number of the eigenvalues $\rho_{i}$ are nonzero. In this basis, the entries $a_{jk}=\langle e_{j},ae_{k}\rangle$ and $a^{\dagger}_{kl}=\langle e_{k},a^{\dagger}e_{l}\rangle$ of the operators $a$ and $a^{\dagger}$ are well defined. Hence, with summation in repeated indices,
$\langle\rho,D\rho\rangle_{\rm HS}$ $=$ $\Big{(}2\rho_{i}\delta_{ij}a_{jk}\rho_{k}\delta_{kl}a^{\dagger}_{li}-a^{\dagger}_{kl}a_{lj}\rho_{j}^{2}(t)\delta_{jk}-a_{jk}a^{\dagger}_{kl}\rho_{l}^{2}(t)\delta_{lj}\Big{)}$ (3.3)
$=$ $\Big{(}2\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{kl}a_{lk}\rho_{k}^{2}-a_{jk}a^{\dagger}_{kj}\rho_{j}^{2}\Big{)}$
$=$ $\Big{(}2\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{ki}a_{ik}\rho_{k}^{2}-a_{ik}a^{\dagger}_{ki}\rho_{i}^{2}\Big{)}$
$=$ $a_{ik}a^{\dagger}_{ki}[2\rho_{i}\rho_{k}-\rho_{k}^{2}-\rho_{i}^{2}]=-|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2}\leq 0$
since $a^{\dagger}_{ji}=\overline{a}_{ij}$. Hence, the nonpositivity of $D$ is proved.
Second, let us check the symmetry of $D$:
$\mathop{\rm tr\,}\nolimits[\rho_{1}D\rho_{2}]=\mathop{\rm tr\,}\nolimits[D\rho_{1})^{{\dagger}}\rho_{2}],\qquad\rho_{1},\rho_{2}\in{\cal D}.$ (3.4)
We have
$\mathop{\rm tr\,}\nolimits[\rho_{1}D\rho_{2}]$ $=$ $\mathop{\rm tr\,}\nolimits[\rho_{1}(a\rho_{2}a^{\dagger}-\frac{1}{2}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{2}a^{{\dagger}}a+a^{{\dagger}}\rho_{2}a-\frac{1}{2}aa^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{2}aa^{{\dagger}})]$
$=$ $\mathop{\rm tr\,}\nolimits[\rho_{1}a\rho_{2}a^{\dagger}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}a^{{\dagger}}a+\rho_{1}a^{{\dagger}}\rho_{2}a-\frac{1}{2}\rho_{1}aa^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}aa^{{\dagger}}]$
$=$ $\mathop{\rm tr\,}\nolimits[a^{\dagger}\rho_{1}a\rho_{2}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}a^{{\dagger}}a\rho_{1}\rho_{2}+a\rho_{1}a^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{1}aa^{{\dagger}}\rho_{2}-\frac{1}{2}aa^{{\dagger}}\rho_{1}\rho_{2}]$
$=$ $\mathop{\rm tr\,}\nolimits[(a^{\dagger}\rho_{1}a-\frac{1}{2}\rho_{1}a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho_{1}+a\rho_{1}a^{{\dagger}}-\frac{1}{2}\rho_{1}aa^{{\dagger}}-\frac{1}{2}aa^{{\dagger}}\rho_{1})\rho_{2}]$
$=$ $\mathop{\rm tr\,}\nolimits[(a^{\dagger}\rho_{1}a-\frac{1}{2}a^{{\dagger}}a\rho_{1}-\frac{1}{2}\rho_{1}a^{{\dagger}}a+a\rho_{1}a^{{\dagger}}-\frac{1}{2}aa^{{\dagger}}\rho_{1}-\frac{1}{2}\rho_{1}aa^{{\dagger}})^{{\dagger}}\rho_{2}]=\mathop{\rm tr\,}\nolimits[(D\rho_{1})^{\dagger}\rho_{2}].$
Hence (3.4) is proved.
ii) The calculations above show that the operator $\Delta$} is neither symmetric nor nonpositive. In particular, the calculations (3.1)–(3.3) give in this case
$\mathop{\rm tr\,}\nolimits[\rho D\rho]=|a_{ik}|^{2}[\rho_{i}\rho_{k}-\rho_{k}^{2}]$ (3.5)
which is not enough for the nonpositivity.
@@ 4 Injectivity
Let us prove Theorem 2.3 iii). We must show that $\rho=0$ if $D\rho=0$ and $\rho\in{\cal D}$. In the case $D\rho=0$, we have also $\langle\rho,D\rho\rangle=0$. The formula (3.3) gives
$\langle\rho,D\rho\rangle_{\rm HS}=-|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2},\qquad\rho\in{\cal D}.$ (4.1)
It implies that $\langle\rho,D\rho\rangle=0$ only if $a_{ik}=0$ for $\rho_{i}\neq\rho_{k}$. Therefore, the eigenspaces corresponding to the eigenvalues $\rho_{i}$ are invariant with respect to $a$ and $a^{\dagger}$. However, there is only one such eigenspace, the entire Hilbert space $X$. Accordingly, all eigenvalues $\rho_{i}$ coinside. Hence, all the eigenvalues are zeros since the zero is in spectrum of any density operator $\rho\in{\cal D}$.
@@
The last step uses the fact that $\rho\in{\cal D}$. Alternatively, we can use the argument that for $\rho\in{\cal D}$ with identical nonzero eigenvalues, the Hilbert–Schmidt norm (2.4) is infinite. This argument also works for general $\rho\in{\rm HS}$ if the relation (4.1) holds. However, our proof of this relation is correct only for $\rho\in{\cal D}$.
@@ 5 Acknowledgements
The authors thank S. Kuksin, M.I. Petelin, A. Shnirelman and H. Spohn for longterm fruitful discussions, and the Institute of Mathematics of BOKU University for the support and hospitality. The research is supported by Austrian Science Fund (FWF) PAT 3476224.
6 Conflict of interest
We have no conflict of interest.
7 Data availability statement
The manuscript has no associated data.

### v4

@@ Abstract
We nonpositivity of two basic dissipation operators is proved in the framework of the theory of the Hilbert space of the Hilbert-Schmidt operators. We show that one of the operators is symmetric, while the other is not.
@@ 2 Dissipation operators and main results
$\rho=\sum_{n,n^{\prime}}\,\sum_{s,s^{\prime}=s_{\pm}}\rho_{n,s;n^{\prime},s^{\prime}}|n,s\rangle\otimes\langle n^{\prime},s^{\prime}|.$ (2.3)
@@
$\rho_{n,s;n^{\prime},s^{\prime}}=\langle n,s|\rho|n^{\prime},s^{\prime}\rangle,$ (2.4)
@@
$\|\rho\|_{\rm HS}^{2}=\mathop{\rm tr\,}\nolimits[\rho^{2}]=\sum_{n,n^{\prime}=0}^{\infty}\,\sum_{s,s^{\prime}=s_{\pm}}|\rho_{n,s;n^{\prime},s^{\prime}}|^{2}<\infty.$ (2.5)
@@
i) Both dissipation operators $D$ and $\Delta$} are defined on the domain ${\cal D}\subset{\rm HS}$ and nonpositive:
$\langle\rho,D\rho\rangle_{\rm HS}\leq 0,\qquad\langle\rho,{\Delta}\rho\rangle_{\rm HS}\leq 0,\qquad\rho\in{\cal D}.$ (2.6)
ii) The operator $D:{\cal D}\to{\cal D}$ is symmetric:
$\langle\rho_{1},D\rho_{2}\rangle_{\rm HS}=\langle D\rho_{1},\rho_{2}\rangle_{\rm HS},\qquad\rho_{1},\rho_{2}\in{\cal D},$ (2.7)
while $\Delta$} is not symmetric.
iii) Both operators $D,{\Delta}:{\cal D}\to{\cal D}$ are injective.
@@ 3 Nonpositivity and symmetry
Here we prove Theorem 2.3 i) and ii). First, let us prove the nonpositivity of $\Delta$}. For $\rho\in{\cal D}$,
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}$ $=$ $\mathop{\rm tr\,}\nolimits\big(\rho{\Delta}\rho\big)=\mathop{\rm tr\,}\nolimits\Big(\rho\big(a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a\big)\Big)$ (3.1)
$=$ $\mathop{\rm tr\,}\nolimits\big(\rho a\rho a^{\dagger}-\rho a^{\dagger}a\rho\big)=\mathop{\rm tr\,}\nolimits\big(\rho a\rho a^{\dagger}-a^{\dagger}a\rho^{2}\big).$
@@
Now we use the fact that $\rho$ is a finite rank Hermitian operator (2.3). Then (1.2) implies that the operators $\rho a\rho a^{\dagger}$ and $a^{\dagger}a\rho^{2}$ have only finite number of nonzero entries (2.4), so their traces are well-defined. Moreover, $\rho$ admits a finite spectral resolution in the orthonormal basis of its eigenvectors $e_{i}\in X_{\infty}$:
$\rho=\sum_{i=1}^{\nu}\rho_{i}e_{i}\otimes e_{i}.$ (3.2)
@@
In this basis, the entries $\rho_{ij}=\rho_{i}\delta_{ij}$, and the entries $a_{jk}=\langle e_{j},ae_{k}\rangle$ and $a^{\dagger}_{kl}=\langle e_{k},a^{\dagger}e_{l}\rangle$ of the operators $a$ and $a^{\dagger}$ are well-defined. Hence, (3.1) implies, with summation in repeated indices,
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}$ $=$ $\rho_{i}\delta_{ij}a_{jk}\rho_{k}\delta_{kl}a^{\dagger}_{li}-a^{\dagger}_{kl}a_{lj}\rho_{j}^{2}\delta_{jk}=\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{kl}a_{lk}\rho_{k}^{2}$ (3.3)
$=$ $\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{ki}a_{ik}\rho_{k}^{2}=a_{ik}a^{\dagger}_{ki}(\rho_{i}\rho_{k}-\rho_{k}^{2})$
$=$ $\frac{1}{2}\Big(a_{ik}a^{\dagger}_{ki}(\rho_{i}\rho_{k}-\rho_{k}^{2})+a_{ki}a^{\dagger}_{ik}(\rho_{k}\rho_{i}-\rho_{i}^{2}))=-\frac{1}{2}|a_{ik}|^{2}(\rho_{i}-\rho_{k})^{2}\leq 0$
since $a^{\dagger}_{ik}=\overline{a}_{ki}$. Now, the nonpositivity is proved for $\Delta$}.
Lemma 3.1.
$D={\Delta}+{\Delta}^{\dagger}$, where $\Delta$$^{\dagger}$ differs from $\Delta$} by swapping $a$ and $a^{\dagger}$:
${\Delta}^{\dagger}\rho=a^{\dagger}\rho a-\frac{1}{2}\rho a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho,\qquad\rho\in{\cal D}.$ (3.4)
Proof.
For $\rho_{1},\rho_{2}\in{\cal D}$,
$\mathop{\rm tr\,}\nolimits[\rho_{1}({\Delta}\rho_{2})]$ $=$ $\mathop{\rm tr\,}\nolimits\Big(\rho_{1}(a\rho_{2}a^{\dagger}-\frac{1}{2}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{2}a^{{\dagger}}a)\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big(\rho_{1}a\rho_{2}a^{\dagger}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}a^{{\dagger}}a\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big(a^{\dagger}\rho_{1}a\rho_{2}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}a^{{\dagger}}a\rho_{1}\rho_{2}\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big((a^{\dagger}\rho_{1}a-\frac{1}{2}\rho_{1}a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho_{1})\rho_{2}\Big)=\mathop{\rm tr\,}\nolimits[({\Delta}^{\dagger}\rho_{1})\rho_{2}].$
Hence, (3.4) is proved. ∎
As the corollary, the nonpositivity of $\Delta$$^{\dagger}$ holds by the same arguments as for $\Delta$} just by swapping $a$ and $a^{\dagger}$, and hence, $D={\Delta}+{\Delta}^{\dagger}$ is also nonpositive. Moreover, $D$ is symmetric on $\cal D$}, while $\Delta$} is not.
Remark 3.2.
The proof of the nonpositivity essentially depends on the symmetry of $\rho$.
@@ 4 Injectivity
Let us prove Theorem 2.3 iii). The formula (3.3) gives
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}=-\frac{1}{2}|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2},\qquad\rho\in{\cal D}.$ (4.1)
It implies that $\langle\rho,{\Delta}\rho\rangle=0$ only if $a_{ik}=0$ for $\rho_{i}\neq\rho_{k}$. Therefore, the eigenspaces of $\rho$ corresponding to the eigenvalues $\rho_{i}$ are invariant with respect to $a$ and $a^{\dagger}$. However, there is only one such eigenspace, the entire Hilbert space $X$. Accordingly, all eigenvalues $\rho_{i}$ coinside. Hence, all the eigenvalues are zeros since the zero is in spectrum of any density operator $\rho\in{\cal D}$.
@@
The last step uses the fact that $\rho\in{\cal D}$. Alternatively, we can use the argument that the Hilbert–Schmidt norm (2.5) is infinite if all the eigenvalues $\rho_{i}$ are nonzero and identical. This argument also works for general $\rho\in{\rm HS}$ if the relation (4.1) holds. However, our proof of this relation is correct only for $\rho\in{\cal D}$.
@@ 5 Acknowledgements
The authors thank S. Kuksin, M.I. Petelin, A. Shnirelman and H. Spohn for longterm fruitful discussions.